You measure 28 turtles' weights, and find they have a mean weight of 64 ounces. Assume the population standard deviation is 11.9 ounces. Based on this, construct a 90% confidence interval for the true population mean turtle weight. Give your answers as decimals, to two places.
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 64

Population standard deviation =    = 11.9

Sample size = n =28

At 90% confidence level the z is ,

= 1 - 90% = 1 - 0.90 = 0.1

/ 2 = 0.1 / 2 = 0.05

Z/2 = Z0.05 = 1.645 ( Using z table )


Margin of error = E = Z/2 * ( /n)

= 1.645* ( 11.9/   28)

= 3.70
At 90% confidence interval estimate of the population mean
is,

- E < < + E

64- 3.70 <   < 64+ 3.70

60.30 <   < 67.70

(60.30 , 67.70)